# in Post-Training and Test-Time Scaling of Large Language Models

### v1

@@ 3.4 Training with Rule-based Rewards
Recently, training with rule-based rewards  [^1] While rule-based rewards may overlap with previous scalar or critique rewards, here we present them separately due to their emerging significance. has gained prominence, driven by the success of DeepSeek-R1 DeepSeek-AI et al. (2025). Rule-based rewards are derived by verifying outputs against specific rules, such as format constraints and evaluation metrics. By leveraging reinforcement learning with rule-based rewards, DeepSeek-R1 demonstrates that language models can acquire long Chain-of-Thoughts (long CoT) abilities for test-time scaling. This enables it to solve complex tasks such as mathematics and coding and show self-reflection and self-correction behaviors. This phenomenon, characterized by a sudden emergence of advanced reasoning capabilities, is referred to as RL scaling or the “Aha moment". Note that the literature sometimes refers to rule-based rewards as verifiable rewards/outcomes due to their clean evaluation criteria.
@@
Later, many following studies have been proposed. DAPO Yu et al. (2025b) and Open-R1 Face (2025) introduce open-source training frameworks, and some extended GRPO algorithms are introduced Xu et al. (2025b); Zuo et al. (2025); Feng et al. (2025c); Zhang et al. (2025b). Logic-RL Xie et al. (2025a) combines accuracy and format rewards to improve logical reasoning through REINFORCE++ Hu (2025). Visual-RFT Liu et al. (2025h) introduces a set of rule-based rewards for visual tasks, such as Intersection over Union (IoU), confidence-based scoring, and format compliance to fine-tune a vision language model via GRPO. These visual tasks include image classification, reasoning grounding, and object detection. Similarly, CLS-RL Li et al. (2025c) designs rule-based rewards for image classification. R1-VL Zhang et al. (2025a) proposes StepGRPO that extends GRPO into step-level multimodal reasoning.
@@ 3.5 Training with Process Rewards
Alternatively, other methods estimate process rewards without explicit annotations. Math-Shepherd Wang et al. (2023b) applies Monte Carlo estimation to infer step-level scores of mathematical reasoning based on the likelihood of reaching a correct final answer. It uses these estimated annotations to train a PRM and then fine-tunes the language model through step-by-step PPO. Jiao et al. (2024b) follow the same way but use DPO training with sampled reasoning trajectories as preference data. PQM Li and Li (2024) reformulates the PRM training as a Q-value ranking problem. It trains a PRM through a margin-based ranking loss over step pairs, encouraging higher Q-values for steps leading to correct answers. OmegaPRM Luo et al. (2024) adopts an efficient and fully automated method to collect step-level supervision with a novel MCTS algorithm. This significantly reduces the annotation cost and enables to scale up the training of high-quality process reward models. HRM Wang et al. (2025b) also leverages reasoning trajectories from MCTS for training. It evaluates both individual steps and consecutive step pairs to capture multi-step coherence and error recovery. Thus this avoids penalizing an early error regardless of subsequent correction.
@@ 4.1 Generate-then-Rank
As aforementioned, outcome reward models may struggle to discern the nuance among candidate responses. Thus many methods adopt process reward models (PRMs) for the generate-then-rank strategy. These methods score intermediate steps of candidate responses through a PRM and aggregate these step-level scores through multiplication or minimum to compute an overall score for ranking or voting Zhang et al. (2025g).
@@ 5.2 Correction with External Feedback
§ 6.2, fill=benchmark-fill, draw=benchmark-border [ MathCheck-GSM Zhou et al. (2024d); MR-GSM8K Zeng et al. (2023);
@@ 6 Benchmarking Reward Models
As reward models play a central role in the learning-from-rewards paradigm for post-training and test-time scaling, rigorous and diverse benchmarks are essential for evaluating their capabilities, like other benchmarks for LLMs Pan et al. (2024b, 2023c); Zhou et al. (2024b); Wu et al. (2022, 2023a, 2023b, 2024c, 2024d, 2024e, 2024b, 2024f, 2024g); Wu (2024). As illustrated in Figure 10, recent benchmarking efforts primarily rely on expert human annotators or AI annotators (e.g., LLM-as-a-Judge frameworks) followed by human verification to ensure reliability. The resulting annotations are mainly pointwise (e.g., scalar scoring) or pairwise (e.g., selecting the preferred response given two options). These benchmarks vary in task coverage, evaluation protocols, annotation sources, and reward formats. We discuss the representative benchmarks in the literature as follows.
@@ 6.1 Benchmarking Outcome Reward Models
A dominant line of benchmarking studies centers on outcome reward models that evaluate the overall quality of generated outputs. Zheng et al. (2023) is an early work that evaluates LLMs’ judging ability by directly prompting them. As LLMs can naturally function as generative reward models, this study also represents one of the earliest benchmarks for reward models. RewardBench Lambert et al. (2024) is the first comprehensive benchmarks for reward models. It aggregates preference data from existing datasets, such as AlpacaEval and MTBench, to evaluate reward model performance in chatting, reasoning, and safety. RM-Bench Liu et al. (2024d) introduces evaluation for reward models on sensitivity to subtle content changes and robustness to style biases. It constructs preference pairs across chat, code, math, and safety domains using GPT-4o. AceMath-RewardBench Liu et al. (2024e) focuses on math-specific evaluations. It tests whether reward models can identify correct solutions from candidates across various mathematical tasks and difficulty levels. RMB Zhou et al. (2024a) furthermore broadens the evaluation scope to 49 real-world scenarios.
@@ 6.2 Benchmarking Process Reward Models
Recently more benchmarks focus on process reward models due to their increasing significance. In detail, several benchmarks focus on math reasoning, such as MathCheck-GSM Zhou et al. (2024d), MR-GSM8K Zeng et al. (2023), and MR-MATH Xia et al. (2024). They require reward models to locate the first error step in a math reasoning solution. Their testing samples are adapted from existing math datasets, including GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). Furthermore, ProcessBench Zheng et al. (2024) features diversity and higher difficulty levels by scaling this up to Olympiad- and competition-level math problems He et al. (2024); Gao et al. (2024b). Beyond step correctness, PRMBench Song et al. (2025b) offers a more fine-grained benchmark. It annotates each step in the reasoning path with specific error types grouped into three dimensions: simplicity, soundness, and sensitivity. The annotations come from LLM-generated perturbations and are subsequently verified by human annotators.
@@ 7.4 Multimodal Tasks
Learning-from-rewards strategies have been widely applied to multimodal tasks, including multimodal understanding and generation. Most studies adopt reinforcement learning and reward-guided decoding methods. For instance, Q-Insight Li et al. (2025d) focuses on improving comprehensive image quality understanding with reinforcement learning. VLM-R1 Shen et al. (2025a) applies reinforcement learning to fine-tune vision-language models and focuses on two tasks: referring expression compression and object detection. Vision-R1 Huang et al. (2025) enhances multimodal reasoning of vision-language models for mathematical VQA. Zhan et al. (2025) proposes another Vision-R1, but it mainly facilitates object localization tasks with vision-language models.
Video-R1 Feng et al. (2025b), VideoChat-R1 Li et al. (2025f), and TinyLLaVA-Video-R1 Zhang et al. (2025e) apply GRPO into video reasoning. R1-V Chen et al. (2025a) and CrowdVLM-R1 Wang et al. (2025e) focus on visual counting. More example applications include multimodal reasoning Zhou et al. (2025b); Meng et al. (2025); Tan et al. (2025); Li et al. (2025b); Liu et al. (2025f), object detection Liu et al. (2025h), segmentation Liu et al. (2025d), and image/video generation Guo et al. (2025b); Liu et al. (2025a).
@@ 7.5 Agents
LLM Agent is an autonomous system that automatically performs complex tasks through task decomposition and action execution in dynamic environments Wang et al. (2024c). Various learning-from-rewards strategies have been applied to training or guiding the agents. AgentRM Xia et al. (2025) targets general-purpose decision-making agents across domains such as web navigation, embodied planning, text games, and tool use. During inference, a reward model guides the agents to choose candidate actions or trajectories. AgentPRM Choudhury (2025) trains LLM agents with a process reward model. KBQA-o1 Luo et al. (2025) guides MCTS with a reward model for the knowledge base question answering task with agents. DeepResearch OpenAI (2025) and DeepResearcher Zheng et al. (2025) design agents for research tasks. They both use reinforcement learning to fine-tune the agents. UI-R1 Lu et al. (2025) introduces a rule-based reinforcement learning framework for GUI action prediction with multimodal agents. InfiGUI-R1 Liu et al. (2025c) is a similar work with GUI agents. RAGEN Wang et al. (2025f) propose training the agents via multi-turn reinforcement learning with a new algorithm based on GRPO.
@@ 7.6 Other Applications
Embodied AI is essential for the development of artificial general intelligence. AI systems, such as embodied robots, must interact with the physical world and complete complex tasks through high-level planning and low-level control. They generally aim to enhance the embodied reasoning abilities with reinforcement learning, such as Cosmos-reason1 Azzolini et al. (2025), iRe-VLA Guo et al. (2025a), Embodied-Reasoner Zhang et al. (2025d), and Embodied-R Zhao et al. (2025a).
@@ 8 Challenges and Future Directions
In this section, we discuss the current challenges and future directions of learning from rewards. Figure 12 summarizes the key challenges and future directions from the perspective of reward model design and learning strategies. Ultimately, we envision the development of interpretable, robust, and continually evolving agent systems capable of interacting with and adapting to the complexities of the real world.
@@ 8.3 Reward Hacking
Reward hacking is a fundamental challenge in learning from rewards Everitt et al. (2021); Amodei et al. (2016); Weng (2024); Liu et al. (2025b). It occurs when models exploit unintended shortcuts in the reward function to obtain high rewards without truly learning the desired behaviors or completing the task as designed. This phenomenon has been observed across domains. For instance, LLMs may fabricate plausible yet incorrect answers, and code LLMs subtly modify unit tests to pass evaluations Denison et al. (2024). Reward backing can also happen during inference, called in-context reward hacking Pan et al. (2024c, a). It arises in self-refinement loops where the same model acts as both the generator and the judge. In such cases, the model may learn to produce outputs that exploit its own evaluation heuristics, leading to inflated internal scores while deviating from true objectives.
@@ 9 Conclusion
In this paper, we present a comprehensive survey of learning LLMs from rewards. We categorize the landscape into three key stages-training, inference, and post-inference, each reflecting a distinct paradigm for integrating reward signals into steering LLMs’ behavior. For each stage, we review representative strategies in terms of reward signal design and learning algorithms. In addition, we summarize recent progress in benchmarking reward models and applications. Finally we identify core challenges and outline promising future directions. We hope this survey provides a structured understanding of the field and inspires future research.

### v2

@@ 3.5 Training with Rule-based Rewards
Recently, training with rule-based rewards  [^2] While rule-based rewards may overlap with previous scalar or critique rewards, here we present them separately due to their emerging significance. has gained prominence due to the success of DeepSeek-R1 DeepSeek-AI et al. (2025). Rule-based rewards are derived by verifying outputs against specific rules, such as format constraints and evaluation metrics. By leveraging reinforcement learning with rule-based rewards, DeepSeek-R1 demonstrates that language models can acquire long Chain-of-Thoughts (long CoT) abilities for test-time scaling. This enables it to solve complex tasks such as mathematics and coding and show self-reflection and self-correction behaviors. This phenomenon, characterized by a sudden emergence of advanced reasoning capabilities, is referred to as RL scaling or the “Aha moment". Note that the literature sometimes refers to rule-based rewards as verifiable rewards/outcomes due to their clean evaluation criteria, for example Reinforcement Learning with Verifiable Rewards (RLVR).
@@
Later, many following studies have been proposed. DAPO Yu et al. (2025b) and Open-R1 Face (2025) introduce open-source training frameworks, and some extended GRPO algorithms are introduced Xu et al. (2025c); Zuo et al. (2025); Feng et al. (2025c); Zhang et al. (2025b). Logic-RL Xie et al. (2025a) combines accuracy and format rewards to improve logical reasoning through REINFORCE++ Hu (2025). Visual-RFT Liu et al. (2025h) introduces a set of rule-based rewards for visual tasks, such as Intersection over Union (IoU), confidence-based scoring, and format compliance to fine-tune a vision language model via GRPO. These visual tasks include image classification, reasoning grounding, and object detection. Similarly, CLS-RL Li et al. (2025c) designs rule-based rewards for image classification. R1-VL Zhang et al. (2025a) proposes StepGRPO that extends GRPO into step-level multimodal reasoning.
Recently, Shao et al. (2025) challenges reinforcement learning with rule-based rewards: they find that random and incorrect rewards can work as correct rewards for some LLMs like QWen2.5-Math.
@@ 3.6 Training with Process Rewards
Alternatively, other methods estimate process rewards without explicit annotations. Math-Shepherd Wang et al. (2023b) applies Monte Carlo estimation to infer step-level scores of mathematical reasoning based on the likelihood of reaching a correct final answer. It uses these estimated annotations to train a PRM and then fine-tunes the language model through step-by-step PPO. Jiao et al. (2024b) follow the same way but use DPO training with sampled reasoning trajectories as preference data. PQM Li and Li (2024) reformulates the PRM training as a Q-value ranking problem. It trains a PRM through a margin-based ranking loss over step pairs, encouraging higher Q-values for steps leading to correct answers. OmegaPRM Luo et al. (2024) adopts an efficient and fully automated method to collect step-level supervision with a novel MCTS algorithm. This significantly reduces the annotation cost and enables to scale up the training of high-quality process reward models. HRM Wang et al. (2025b) also leverages reasoning trajectories from MCTS for training. It evaluates both individual steps and consecutive step pairs to capture multi-step coherence and error recovery. Thus this avoids penalizing an early error regardless of subsequent correction. SCOPE Xu et al. (2025b) compresses math reasoning steps into syntax trees, which avoids large-scale sampling and thus reduces annotation cost.
@@ 4.1 Generate-then-Rank
As aforementioned, outcome reward models may struggle to discern the nuance among candidate responses. Thus many methods adopt process reward models (PRMs) for the generate-then-rank strategy. These methods score intermediate steps of candidate responses through a PRM and aggregate these step-level scores through multiplication or minimum to compute an overall score for ranking or voting Zhang et al. (2025h).
@@ 5.2 Correction with External Feedback
§ 6.2, fill=benchmark-fill, draw=benchmark-border [ MathCheck-GSM Zhou et al. (2024c); MR-GSM8K Zeng et al. (2023);
@@ 6 Benchmarking Reward Models
As reward models play a central role in the learning-from-rewards paradigm for post-training and test-time scaling, rigorous and diverse benchmarks are essential for evaluating their capabilities. As illustrated in Figure 10, recent benchmarking efforts primarily rely on expert human annotators or AI annotators (e.g., LLM-as-a-Judge frameworks) followed by human verification to ensure reliability. The resulting annotations are mainly pointwise (e.g., scalar scoring) or pairwise (e.g., selecting the preferred response given two options). These benchmarks vary in task coverage, evaluation protocols, annotation sources, and reward formats. We discuss the representative benchmarks in the literature as follows.
@@ 6.1 Benchmarking Outcome Reward Models
A dominant line of benchmarking studies centers on outcome reward models that evaluate the overall quality of generated outputs. Zheng et al. (2023) is an early work that evaluates LLMs’ judging ability by directly prompting them. As LLMs can naturally function as generative reward models, this study also represents one of the earliest benchmarks for reward models. Zhou et al. (2025c) follow a similar idea and focus on evaluation for test-time scaling.
RewardBench Lambert et al. (2024) is the first comprehensive benchmarks for reward models. It aggregates preference data from existing datasets, such as AlpacaEval and MTBench, to evaluate reward model performance in chatting, reasoning, and safety. RM-Bench Liu et al. (2024d) introduces evaluation for reward models on sensitivity to subtle content changes and robustness to style biases. It constructs preference pairs across chat, code, math, and safety domains using GPT-4o. AceMath-RewardBench Liu et al. (2024e) focuses on math-specific evaluations. It tests whether reward models can identify correct solutions from candidates across various mathematical tasks and difficulty levels. RMB Zhou et al. (2024a) furthermore broadens the evaluation scope to 49 real-world scenarios.
@@ 6.2 Benchmarking Process Reward Models
Recently more benchmarks focus on process reward models due to their increasing significance. In detail, several benchmarks focus on math reasoning, such as MathCheck-GSM Zhou et al. (2024c), MR-GSM8K Zeng et al. (2023), and MR-MATH Xia et al. (2024). They require reward models to locate the first error step in a math reasoning solution. Their testing samples are adapted from existing math datasets, including GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). Furthermore, ProcessBench Zheng et al. (2024) features diversity and higher difficulty levels by scaling this up to Olympiad- and competition-level math problems He et al. (2024); Gao et al. (2024b). Beyond step correctness, PRMBench Song et al. (2025b) offers a more fine-grained benchmark. It annotates each step in the reasoning path with specific error types grouped into three dimensions: simplicity, soundness, and sensitivity. The annotations come from LLM-generated perturbations and are subsequently verified by human annotators.
@@ 7.4 Multimodal Tasks
Learning-from-rewards strategies have been widely applied to multimodal tasks, including multimodal understanding and generation. Most studies adopt reinforcement learning and reward-guided decoding methods. For instance, Q-Insight Li et al. (2025d) focuses on improving comprehensive image quality understanding with reinforcement learning. VLM-R1 Shen et al. (2025a) applies reinforcement learning to fine-tune vision-language models and focuses on two tasks: referring expression compression and object detection. Vision-R1 Huang et al. (2025b) enhances multimodal reasoning of vision-language models for mathematical VQA. Zhan et al. (2025) proposes another Vision-R1, but it mainly facilitates object localization tasks with vision-language models.
Video-R1 Feng et al. (2025b), VideoChat-R1 Li et al. (2025f), and TinyLLaVA-Video-R1 Zhang et al. (2025f) apply GRPO into video reasoning. R1-V Chen et al. (2025a) and CrowdVLM-R1 Wang et al. (2025f) focus on visual counting. More example applications include multimodal reasoning Zhou et al. (2025b); Meng et al. (2025); Tan et al. (2025); Li et al. (2025b); Liu et al. (2025f), object detection Liu et al. (2025h), segmentation Liu et al. (2025d), and image/video generation Guo et al. (2025c); Liu et al. (2025a).
@@ 7.5 Agents
LLM Agent is an autonomous system that automatically performs complex tasks through task decomposition and action execution in dynamic environments Wang et al. (2024c). Various learning-from-rewards strategies have been applied to training or guiding the agents. AgentRM Xia et al. (2025) targets general-purpose decision-making agents across domains such as web navigation, embodied planning, text games, and tool use. During inference, a reward model guides the agents to choose candidate actions or trajectories. AgentPRM Choudhury (2025) trains LLM agents with a process reward model. KBQA-o1 Luo et al. (2025) guides MCTS with a reward model for the knowledge base question answering task with agents. DeepResearch OpenAI (2025) and DeepResearcher Zheng et al. (2025) design agents for research tasks. They both use reinforcement learning to fine-tune the agents. UI-R1 Lu et al. (2025) introduces a rule-based reinforcement learning framework for GUI action prediction with multimodal agents. InfiGUI-R1 Liu et al. (2025c) is a similar work with GUI agents. RAGEN Wang et al. (2025g) propose training the agents via multi-turn reinforcement learning with a new algorithm based on GRPO.
@@ 7.6 Other Applications
Embodied AI is essential for the development of artificial general intelligence. AI systems, such as embodied robots, must interact with the physical world and complete complex tasks through high-level planning and low-level control. They generally aim to enhance the embodied reasoning abilities with reinforcement learning, such as Cosmos-reason1 Azzolini et al. (2025), iRe-VLA Guo et al. (2025b), Embodied-Reasoner Zhang et al. (2025d), and Embodied-R Zhao et al. (2025a).
@@ 8 Challenges and Future Directions
In this section, we discuss the current challenges and future directions of learning from rewards. Figure 12 summarizes the key challenges and future directions from the perspective of reward models and learning strategies. Ultimately, we envision the development of interpretable, robust, and continually evolving agent systems capable of interacting with and adapting to the complexities of the real world.
@@ 8.3 Reward Hacking
Reward hacking is a fundamental challenge in learning from rewards Everitt et al. (2021); Amodei et al. (2016); Weng (2024); Liu et al. (2025b). It occurs when models exploit unintended shortcuts in the reward function to obtain high rewards without truly learning the desired behaviors or completing the task as designed. This phenomenon has been observed across domains. For instance, LLMs may fabricate plausible yet incorrect answers, and code LLMs subtly modify unit tests to pass evaluations Denison et al. (2024). Reward backing can also happen during inference, called in-context reward hacking Pan et al. (2024b, a). It arises in self-refinement loops where the same model acts as both the generator and the judge. In such cases, the model may learn to produce outputs that exploit its own evaluation heuristics, leading to inflated internal scores while deviating from true objectives.
@@ 9 Conclusion
In this paper, we present a comprehensive survey of learning from rewards in LLMs. We categorize the landscape into three key stages—training, inference, and post-inference, each reflecting a distinct paradigm for integrating reward signals into steering LLMs’ behavior. For each stage, we review representative studies in terms of reward models and learning strategies. In addition, we summarize recent progress in benchmarking reward models and applications. Finally we identify core challenges and outline promising future directions. We hope this survey provides a structured understanding of the field and inspires future research.